solve this system using substitution . show work
x+y=-4
x-y=2

Answer by mananth(16946)

x+y= -4 .............1
-y=2-x ---------2
Substitute y in (1)
x+(-2+ x)= -4
x-2+ x=-4
x+x=-4
2x = -2
/ 2
x= -1
Plug the value of x in (1)
x+y = -4
1*-1 +y= -4
-1 + y = -4
y = -3
m.ananth@hotmail.ca